# SafeFix: Targeted Model Repair via Controlled Image Generation

### v1

@@ 5.5 Effect of Alternative Bug Slice Selections
# images Base RH BS SE RH_BS RH_SE RH_BS_SE
1,000 90.68 91.31 91.38 91.68 91.30 91.55 91.69
5,000 91.59 91.76 91.68 91.92 91.19 91.38
10,000 91.61 91.54 91.76 91.51 91.45 91.71
Table 3: Validation accuracy (%) for Base and attribute-variant selections. Ours is the RH_BS_SE combination.
Table 3 reports validation accuracy on CelebA using ResNet when augmenting with different subsets of rare-case attributes. RH, BS, and SE refer to red hair, brown skin, and sad emotion, respectively. Combined settings such as RH_BS and RH_BS_SE(Ours) correspond to augmenting images exhibiting multiple rare-case attributes. Across all settings, augmentation improves over the base model (90.68%). Notably, our three-way combination RH_BS_SE achieves the highest accuracy at 1,000 and 10,000 images, and remains competitive at 5,000 images, demonstrating that targeting intersecting rare-case conditions yields more effective repair than augmenting isolated attributes alone. While SE alone achieves the highest accuracy at 10,000 images (91.76%), augmenting only the sad-emotion slice fails to address other concurrent failure modes and can even introduce new biases. Augmenting solely on the SE slice leaves that subset still underperforming in validation (90.66%), indicating persistent hair-color errors. In contrast, our RH_BS_SE combination improves all three slices at once as indicated by Figure 3(a). For the attribute combination RH_BS_SE, the number of bugs drops from 10 to 4, while augmenting only the SE attribute reduces the count to 7.
A similar pattern holds on ImageNet10. Selecting the pink color + fabric texture slice reduces the number of rare-case bugs for ResNet from 13 to 7, whereas augmenting only pink color reduces them to 9 and augmenting only fabric texture reduces them to 10. The number of identified bugs is sensitive to these thresholds. For our analysis on both datasets, we explored a rare threshold $\rho$ (0–0.05) and the accuracy difference threshold $\epsilon$ (0.01–0.03). Using values outside these ranges can increase the number of identified bugs.
@@ 5.6 Ablation Study
Components CelebA Accuracy (%)
@@
a) 90.93 90.87 90.76
b) 91.15 90.73 91.35
c) 91.35 91.32 91.54
d) 91.69 91.38 91.71
Table 4: Ablation on the impact of CDM and LVLM components at different scales. This highlights SafeFix’s robustness under both sparse and saturated augmentations.
We ablate the contributions of the two core components in our pipeline: the Conditional Diffusion Model (CDM) and the LVLM filter. Table 4 reports the validation accuracy on CelebA when using different combinations of these components across three augmentation scales. Note that configuration (a) corresponds to the baseline strategy from HiBug Chen et al. (2023). We observe several trends: adding either the LVLM (b) or the CDM (c) individually improves accuracy across all scales; combining both the CDM and the LVLM (d) yields the highest performance in all cases.
@@ 6 Conclusion
We presented an automated model repair pipeline for vision tasks that combines failure-attribute diagnostics with targeted synthetic augmentation. We use a Conditional Diffusion Model (CDM) to generate attribute-preserving variants and apply LVLM-based filtering to ensure semantic correctness, thereby focusing augmentation on true rare-case failure slices. Experiments on CelebA and ImageNet10 with ResNet, ViT, and CLIP backbones show consistent gains in accuracy and reduced bugs on underrepresented subpopulations, outperforming CDM-based and HiBug baselines. Ablations confirm that the CDM and the LVLM contribute complementary benefits, highlighting the importance of targeted, validated augmentation for robust model repair.
Limitations. Our method’s effectiveness is fundamentally constrained by its components. Crucially, the pipeline can inherit and amplify biases from the diffusion generator or the LVLM filter, potentially perpetuating the fairness issues we aim to address if these components have demographic performance gaps. Other limitations include the significant computational cost of generation and filtering, potential image artifacts, and a fixed attribute vocabulary that cannot address unmodeled failure modes. Future work will focus on mitigating inherited biases, improving efficiency, and dynamic attribute discovery.
@@ Appendix A Attribute Assignment Details
CelebA. Attribute assignments for the CelebA dataset are based on the official annotations provided by the dataset and follow the settings used in HiBug Chen et al. [2023].
@@ Appendix B Conditional Diffusion Model Details
We use Stable Diffusion v1.5 as our text-to-image diffusion backbone and apply ControlNet Zhang et al. [2023] to enable fine-grained attribute control. The ControlNet model is conditioned on soft edge maps extracted from original training images using the HED detector Xie and Tu [2015], which preserves structural consistency during generation.
@@ C.1 LVLM Query Format
For semantic filtering of generated images, we use two large vision–language models (LVLMs), specifically Qwen2.5-VL Bai et al. [2025] and LLaVA-v1.5 Liu et al. [2023], to verify whether each image accurately satisfies its intended attribute-value conditions.
@@ Appendix D Additional Implementation Details
We conduct all experiments on an NVIDIA A100 GPU. We take the CelebA “wearing lipstick” classification task as an example. We evaluate three backbone architectures for the classification task. ResNet-18 He et al. [2016], ViT-B/16 Dosovitskiy et al. [2020], and CLIP (ViT-B/32) Radford et al. [2021] are all initialized with random weights. All vision models are trained using cross-entropy loss. Optimization is performed using Adam Kingma and Ba [2015] with a learning rate of $1\times 10^{-4}$ and weight decay of $1\times 10^{-3}$. All experiments use a batch size of 64 and are trained for 20 epochs.
For rare-case bug discovery and augmentation, we take the CelebA “wearing lipstick” classification task as an example. We first use a computer vision model $f_{\theta}\colon\mathcal{X}\to\mathcal{Y}$ to produce predictions and compute the overall validation accuracy $\mathrm{Acc}(\mathcal{D}_{\mathrm{val}})$. For our experiments on the CelebA dataset, we use 7 distinct hair color attributes. These color categories were initially proposed using ChatGPT (GPT-4 with vision OpenAI [2023]) and subsequently assigned to the images by the BLIP Li et al. [2022] model. This assignment process followed the same setting as in Chen et al. [2023]. Accordingly, we set the rare threshold $\rho=0.05$ (which is less than $1/7\approx 0.143$) and the accuracy difference threshold $\epsilon=0.01$. We found there are 2,484 red-hair images out of the 80,000 training samples, which corresponds to only 3.11% of the data. Their slice accuracy is 0.8901, which is lower than the overall $\mathrm{Acc}(\mathcal{D}_{\mathrm{val}})=0.9068$. Based on these thresholds, we identify the subset of red-hair images that satisfy both conditions and treat them as a candidate rare-case bug slice. For rare-case bug discovery and augmentation, we follow the procedure described above, using validation accuracy and slice support to identify candidate attributes. We set the rarity threshold $\rho=0.05$ and the accuracy difference threshold $\epsilon=0.01$ based on dataset statistics. For synthetic augmentation, we use the ControlNet Zhang et al. [2023] as the conditional diffusion model, which is based on Stable Diffusion 1.5 Rombach et al. [2022] and conditioned on soft HED boundaries Xie and Tu [2015]. The soft HED boundary preserves fine structural details from the input images, making this approach particularly suitable for attribute-preserving edits such as recoloring and stylizing. To filter and validate generated images, we employ the Qwen2.5-VL Bai et al. [2025] as the primary large vision–language model (LVLM), and additionally compare results with LLaVA-v1.5 Liu et al. [2023].
@@
Appendix E Large-Scale Settings in Main Results
In the main experiments, we show that SafeFix improves validation accuracy by augmenting underrepresented slices. To further assess its effectiveness under large-scale augmentation, we consider a scenario where a previously rare-case attribute—red hair—becomes common in training after adding 10,000 generated images (around 12,480 red-hair images out of 90,000 total), while its prevalence in the fixed validation set remains low (621 out of 20,000).
This setup creates an important distinction: although red hair is no longer rare in training ($\sim 14\%$), it remains rare in validation ($\sim 3\%$). Nevertheless, the validation accuracy on the red-hair subset improves significantly—from 90.02% (559/621) to 91.47% (567/621) —demonstrating that SafeFix effectively improves generalization to previously underperforming groups.
@@ Appendix E Large-Scale Settings in Main Results
Our fixed validation set continues to reflect real-world statistics, ensuring that improvements in rare-case performance are meaningful. This practice—augmenting rare cases in training while keeping the test distribution fixed—is consistent with standard methodology in prior work Chen et al. [2023]. These findings confirm that SafeFix improves generalization on rare-case bugs without sacrificing reliability or introducing new biases.
Appendix F Additional Experiments
F.1 Generation Quality Analysis
We compare HiBug and SafeFix using four key metrics—LPIPS diversity Zhang et al. [2018], CLIP consistency Radford et al. [2021], Fréchet Inception Distance (FID) Heusel et al. [2017] against real images, and the KL divergence of deep-feature distributions—using the ImageNet10 dataset with the pink color + fabric texture setting as an illustrative example, as shown in Table 5.
@@ F.1 Generation Quality Analysis
Table 5: Generation-quality metrics for HiBug vs. SafeFix on ImageNet10 (pink-color + fabric-texture). “Real→Gen” indicates the Fréchet distance computed between the real and generated image feature distributions; lower FID means the generated set is closer to the real distribution. Arrows indicate preference: $\uparrow$ higher is better; $\downarrow$ lower is better.
SafeFix achieves substantially lower FID and KL divergence compared to HiBug, indicating that its samples lie much closer to the real-image distribution. Although its LPIPS diversity is moderately reduced and CLIP consistency slightly lower, these modest trade-offs yield more reliable, attribute-faithful examples for downstream repair. In practice, fewer but higher-quality variants help correct rare-case failures without introducing noisy or out-of-distribution artifacts.
F.2 Qualitative Analysis of Attribute Editing
@@
In the CelebA dataset, we primarily target rare-case attributes such as red hair (RH), brown skin (BS), and sad emotion (SE), but our framework is flexible and supports other attribute configurations. Figure 5 shows examples of original and SafeFix-generated images.
In the first example, only yellow hair was added to the original face while all other identity features remained unchanged. The second image reflects a transformation into the red hair and brown skin attribute combination. In the third example, the model aimed to edit the hair color to red, but instead altered the background to a reddish hue without changing the hair itself—highlighting a typical failure of generative models when not filtered.
@@
F.3 Evaluation with Mixed Train–Validation Synthetic Split

### v2

@@ 5.5 Effect of Alternative Bug Slice Selections
# images RH BS SE RH_BS RH_SE Ours
1,000 7.85 8.59 11.77 7.74 10.39 14.32
5,000 10.82 12.62 11.77 14.42 6.57 12.69
10,000 11.03 10.29 12.62 9.97 9.33 14.95
Table 4: Fix Rate (FR %) for attribute-variant selections. Ours is the RH_BS_SE combination.
Table 4 reports FR on CelebA using ResNet augmenting with different subsets of rare-case attributes. RH, BS, and SE refer to red hair, brown skin, and sad emotion, respectively. Combined settings like RH_BS and RH_BS_SE (Ours) augment images with multiple attributes. Augmentation improves over the base model (90.57%) across all settings. Notably, our three-way combination, RH_BS_SE, achieves the highest FR at $1,000$ and $10,000$ images and remains competitive at $5,000$ images, showing that targeting intersecting rare-case conditions is more effective than augmenting isolated attributes alone. Although RH_BS alone achieves the highest FR at $5,000$ images, augmenting only that slice does not address concurrent failure modes, indicating persistent errors. Firstly, our RH_BS_SE combination improves all three slices concurrently (Figure 3(a)). Secondly, for the RH_BS_SE attribute combination, the number of bugs decreases from $10$ to $4$, while augmenting only SE or RH_BS reduces the count to $7$ and $6$ respectively (using $\rho=0.05$ and $\epsilon=0.03$). A similar pattern holds on ImageNet10: selecting the pink color + fabric texture slice reduces rare-case bugs from $13$ to $7$, whereas augmenting only pink color reduces them to $9$ and only fabric texture reduces them to $10$.
After further verification, this reduction does not introduce any new bugs that were not present before, indicating that SafeFix repairs existing failures without adding new bugs on both datasets. The number of identified bugs is sensitive to these thresholds and using different threshold values can change the number of bugs fixed, as shown in Appendix G.5. We note that larger attribute combinations (4 to 7 attributes) also operate correctly but reach lower FR (around 9% on CelebA and 7% on ImageNet10), falling below our selected attribute groups.
5.6 Effect of the LVLM Filter and Human Audit
Error Type Before After
BG 7 2
ATTR 24 3
Table 5: Average error rates (%) before and after LVLM filtering.
We use an LVLM as the attribute filter, and a human-based test yields nearly identical results. In a human audit conducted with 5 AI graduate students on 300 CelebA images, the pass rates are 97% for “red hair”, 99% for “brown skin”, 78% for “sad emotion”, and 98% for the original label, with an average overlap of about 95% with Qwen2.5-VL and 93% with LLaVA-v1.5, showing that the LVLMs closely match human verification. The main errors arise from background color changes (BG) ($7\%$) and the edited image failing to express the target attribute (ATTR) ($24\%$); the LVLM fixes most of them, as shown in Table 5.
@@ 5.7 Ablation Study
Components CelebA Fix Rate (FR %)
@@
a) 6.79 4.88 3.92
b) 7.21 4.77 6.99
c) 8.06 8.27 10.07
d) 14.32 12.09 14.95
Table 6: Ablation on the impact of CDM and LVLM components at different scales, reported as Fix Rate (FR %). This highlights SafeFix’s robustness under both sparse and saturated augmentations.
We ablate the contributions of the two core components in our pipeline: the Conditional Diffusion Model (CDM) and the LVLM filter. Table 6 reports the fix rate on CelebA when using different combinations of these components across three augmentation scales. Note that configuration (a) corresponds to the baseline strategy from HiBug [chen2023hibug]. We observe several trends: adding either the LVLM (b) or the CDM (c) individually improves accuracy across all scales; SafeFix combining both the CDM and the LVLM (d) yields the highest performance in all cases.
@@ 6 Conclusion
We presented an automated model repair pipeline for vision tasks that combines failure‐attribute diagnostics with targeted synthetic augmentation. We use a Conditional Diffusion Model (CDM) to generate attribute-preserving variants and apply LVLM-based filtering to ensure semantic correctness, thereby focusing augmentation on true rare-case failure slices. Experiments on CelebA and ImageNet10 with ResNet, ViT, and CLIP backbones show consistent gains in accuracy and reduced bugs on underrepresented subpopulations, outperforming CDM-based and HiBug baselines. Ablations confirm that the CDM and the LVLM contribute complementary benefits, highlighting the importance of targeted, validated augmentation for robust model repair.
Limitations. Our method’s effectiveness is constrained by its components. Crucially, the pipeline can inherit biases from the diffusion model or the LVLM, potentially perpetuating the fairness issues if these components have demographic bias. Other limitations include the computational cost of generation and filtering, potential image artifacts, and a fixed attribute vocabulary that cannot address unmodeled failure modes. Future work will focus on mitigating inherited biases, improving efficiency, and dynamic attribute discovery.
@@ Appendix A Attribute Assignment Details
CelebA. Attribute assignments for the CelebA dataset are based on the official annotations provided by the dataset and follow the settings used in HiBug [chen2023hibug].
@@ Appendix B Conditional Diffusion Model Details
We use Stable Diffusion v1.5 as our text-to-image diffusion backbone and apply ControlNet [zhang2023adding] to enable fine-grained attribute control. The ControlNet model is conditioned on soft edge maps extracted from original training images using the HED detector [xie2015holistically], which preserves structural consistency during generation.
@@ C.1 LVLM Query Format
For semantic filtering of generated images, we use two large vision–language models (LVLMs), specifically Qwen2.5-VL [bai2025qwen25vl] and LLaVA-v1.5 [liu2023improvedllava], to verify whether each image accurately satisfies its intended attribute-value conditions.
@@ Appendix D Additional Implementation Details
We conduct all experiments on an NVIDIA A100 GPU. We take the CelebA “wearing lipstick” classification task as an example. We evaluate three backbone architectures for the classification task. ResNet-18 [he2016deep], ViT-B/16 [dosovitskiy2020vit], and CLIP (ViT-B/32) [radford2021learning] are all initialized with random weights. All vision models are trained using cross-entropy loss. Optimization is performed using Adam [kingma2015adam] with a learning rate of $1\times 10^{-4}$ and weight decay of $1\times 10^{-3}$. All experiments use a batch size of 64 and are trained for 20 epochs.
For rare-case bug discovery and augmentation, we take the CelebA “wearing lipstick” classification task as an example. We first use a computer vision model $f_{\theta}\colon\mathcal{X}\to\mathcal{Y}$ to produce predictions and compute the overall test accuracy $\mathrm{Acc}(\mathcal{D}_{\mathrm{val}})$. For our experiments on the CelebA dataset, we use 7 distinct hair color attributes. These color categories were initially proposed using ChatGPT (GPT-4 with vision [openai2023gpt4]) and subsequently assigned to the images by the BLIP [li2022blip] model. This assignment process followed the same setting as in [chen2023hibug]. Accordingly, we set the rare threshold $\rho=0.05$ (which is less than $1/7\approx 0.143$) and the accuracy difference threshold $\epsilon=0.03$. We found there are 2,484 red-hair images out of the 80,000 training samples, which corresponds to only 3.11% of the data. Their slice accuracy is 0.8901, which is lower than the overall $\mathrm{Acc}(\mathcal{D}_{\mathrm{val}})=0.9068$. Based on these thresholds, we identify the subset of red-hair images that satisfy both conditions and treat them as a candidate rare-case bug slice. For rare-case bug discovery and augmentation, we follow the procedure described above, using accuracy and slice support to identify candidate attributes. We set the rarity threshold $\rho=0.05$ and the accuracy difference threshold $\epsilon=0.03$ based on dataset statistics. For synthetic augmentation, we use the ControlNet [zhang2023adding] as the conditional diffusion model, which is based on Stable Diffusion 1.5 [rombach2022high] and conditioned on soft HED boundaries [xie2015holistically]. The soft HED boundary preserves fine structural details from the input images, making this approach particularly suitable for attribute-preserving edits such as recoloring and stylizing. To filter and validate generated images, we employ the Qwen2.5-VL [bai2025qwen25vl] as the primary large vision–language model (LVLM), and additionally compare results with LLaVA-v1.5 [liu2023improvedllava].
@@
Appendix E Accuracy Results
In the Results section, we introduce the Fix Rate (FR), which can be inferred from accuracy metrics. Therefore, we include the accuracy results in Table 7 and Table 8 for the CelebA and ImageNet10 datasets, respectively.
On CelebA, we select rare-case bugs defined by attribute–value combinations red hair, brown skin, and sad emotion for ResNet and ViT, and yellow hair, brown skin, and sad emotion for CLIP, based on the most frequent patterns identified among failure slices. SafeFix consistently achieves the highest test accuracy across all models (ResNet, ViT, and CLIP) and varying levels of synthetic augmentation as shown in Table 7. For example, with 1,000 added images, our method improves accuracy by +1.35% (ResNet), +2.89% (ViT), and +2.78% (CLIP) relative to their respective baselines. These results show that our attribute-targeted augmentation and filtering pipeline is effective in repairing rare-case failure slices, outperforming both CDM-based and HiBug baselines.
On ImageNet10, similar trends emerge, as shown in Table 8. For ResNet and ViT, we target rare-case bugs involving pink color and fabric texture, while for CLIP we use orange color and fabric texture. Across all models, our proposed method consistently surpasses the baseline methods. Specifically, our method improves ResNet accuracy by +3.18% (at 100 images) compared to the base model. ViT and CLIP also exhibit a steady improvement compared to other methods.
Method ResNet ViT CLIP
Base 90.57 85.02 88.32
DA-CDM 90.95 90.88 90.86 85.63 86.55 86.04 90.11 90.25 90.14
Mask-ControlNet 91.08 91.26 91.15 86.71 86.59 86.95 90.20 90.18 90.28
HiBug_Class 91.05 91.12 90.89 85.79 86.18 86.10 90.03 90.07 89.95
HiBug_Task 91.21 91.03 90.94 87.82 86.78 86.18 90.16 90.09 90.03
Ours (L) 91.55 91.65 91.67 87.73 87.78 87.32 90.92 90.95 90.86
Ours (Q) 91.92 91.71 91.98 87.91 87.33 87.41 90.98 91.02 90.71
Table 7: Test accuracy (%) on CelebA for varying numbers of added images, models, and methods. Ours (L) and Ours (Q) denote using LLaVA-v1.5-7B and Qwen2.5-VL-7B as the large vision-language model filter, respectively.
Method ResNet ViT CLIP
Base 71.73 97.42 93.78
DA-CDM 73.21 73.80 74.14 97.58 97.76 97.77 93.96 94.05 94.15
Mask-ControlNet 73.22 72.54 72.39 97.25 97.69 97.52 93.44 93.88 93.70
HiBug_Class 72.24 72.15 71.94 97.12 97.07 97.28 93.92 94.21 93.83
HiBug_Task 73.51 72.88 73.36 97.66 97.78 97.58 93.29 94.18 94.15
Ours (L) 74.10 74.19 73.92 98.21 98.17 98.11 94.49 94.80 94.44
Ours (Q) 74.31 73.43 74.88 98.22 98.09 98.41 94.57 94.88 94.95
Table 8: Test accuracy (%) on ImageNet10 under varying image counts, models, and methods.
Appendix F Large-Scale Settings in Main Results
In the main experiments, we show that SafeFix improves test accuracy by augmenting underrepresented slices. To further assess its effectiveness under large-scale augmentation, we consider a scenario where a previously rare-case attribute—red hair—becomes common in training after adding 10,000 generated images (around 12,480 red-hair images out of 90,000 total), while its prevalence in the fixed validation set remains low (621 out of 20,000).
This setup creates an important distinction: although red hair is no longer rare in training ($\sim 14\%$), it remains rare in validation ($\sim 3\%$). Nevertheless, the accuracy on the red-hair subset improves significantly, demonstrating that SafeFix effectively improves generalization to previously underperforming groups.
@@ Appendix F Large-Scale Settings in Main Results
Our fixed validation set continues to reflect real-world statistics, ensuring that improvements in rare-case performance are meaningful. This practice—augmenting rare cases in training while keeping the test distribution fixed—is consistent with standard methodology in prior work [chen2023hibug]. These findings confirm that SafeFix improves generalization on rare-case bugs without sacrificing reliability or introducing new biases.
Appendix G Additional Experiments
G.1 Generation Quality Analysis
We compare HiBug and SafeFix using four key metrics—LPIPS diversity [zhang2018unreasonable], CLIP consistency [radford2021learning], Fréchet Inception Distance (FID) [heusel2017gans] against real images, and the KL divergence of deep-feature distributions—using the ImageNet10 dataset with the pink color + fabric texture setting as an illustrative example, as shown in Table 9.
@@ G.1 Generation Quality Analysis
Table 9: Generation‐quality metrics for HiBug vs. SafeFix on ImageNet10 (pink‐color + fabric‐texture). “Real→Gen” indicates the Fréchet distance computed between the real and generated image feature distributions; lower FID means the generated set is closer to the real distribution. Arrows indicate preference: $\uparrow$ higher is better; $\downarrow$ lower is better.
SafeFix achieves substantially lower FID and KL divergence compared to HiBug, indicating that its samples lie much closer to the real‐image distribution. Although its LPIPS diversity is moderately reduced and CLIP consistency slightly lower, these modest trade‐offs yield more reliable, attribute‐faithful examples for downstream repair. In practice, fewer but higher‐quality variants help correct rare-case failures without introducing noisy or out‐of‐distribution artifacts.
G.2 SafeFix with alternative conditional diffusion models
The targeted generation stage in our pipeline uses ControlNet[zhang2023adding] as the conditional diffusion model. Although ControlNet is effective due to its stable structure-preserving behavior, relying on a single generator raises questions about whether SafeFix depends on ControlNet-specific features. To test the generality of the pipeline, we replaced ControlNet with another conditional diffusion model, InstructPix2Pix [brooks2023instructpix2pix], while keeping the LVLM filter (Qwen2.5-VL). The resulting accuracy on CelebA, shown in Table 10, exceeds all baselines reported in the main results and is only slightly lower than SafeFix with ControlNet. This demonstrates that the pipeline is not tied to ControlNet and functions well with alternative generators.
Method ResNet (base acc: 90.57%)
1k Images 5k Images 10k Images
DA-CDM 4.03 3.29 3.08
Mask-ControlNet 5.41 7.32 6.15
HiBug_Class 5.09 5.83 3.40
HiBug_Task 6.79 4.88 3.92
Ours (InstructPix2Pix) 11.32 11.85 11.74
Ours (ControlNet) 14.32 12.09 14.95
Table 10: Fix Rate (FR%) using different conditional diffusion models in the targeted generation stage. InstructPix2Pix [brooks2023instructpix2pix] serves as an alternative generator and shows strong repair performance.
G.3 Qualitative Analysis of Attribute Editing and Failure Cases
Figure 2 shows some good examples generated by the conditional diffusion model, but conditional diffusion models can make mistakes. In the CelebA dataset, we primarily target rare-case attributes such as red hair (RH), brown skin (BS), and sad emotion (SE), but our framework is flexible and supports other attribute configurations. Figure 5 shows examples of original and SafeFix-generated images. In the first example, only yellow hair was added to the original face while all other identity features remained unchanged. The second image reflects a transformation into the red hair and brown skin attribute combination. In the third example, the model aimed to edit the hair color to red, but instead altered the background to a reddish hue without changing the hair itself—highlighting a typical failure of generative models when not filtered.
@@ G.3 Qualitative Analysis of Attribute Editing and Failure Cases
G.4 Evaluating Synthetic Data as a Standalone Training Set